# Data-Driven Compute Overlays for Interactive Geographic Simulation and Visualization

### v1

@@ 1 Background: (Web-Based) GeoVis
Common geographic visualization techniques are glyphs placed on top of maps and color overlays [12]. The focus of our work lies on overlays. Here, we review three principle paradigms to data-driven map overlays particularly for the web:
@@
Map tiles are the foundation for all popular online map applications, where map clients request pre-rasterized map images depending on the currently shown zoom level and visible field of view [13] from a dedicated server – also known as “slippy maps” [8]. Data-driven overlays can be similarly pre-rendered into offline tiles and overlaid onto a base map [6, 15]. A well-known open-source library to build layered maps is OpenLayers [16]. Offline tiles are useful when the data to be overlaid itself is big and static. The disadvantage is that the server has to re-compute the tiles whenever the underlying data changes. Users have no or only very limited possibilities to interact with the data.
@@
An alternative approach is to render the final images to be viewed by the client directly on the server. The client only forwards users’ interaction with the map (e.g., panning and zooming) and with the data (e.g., filtering) to the server and finally shows the image transmitted by the server. The application is therefore largely independent from the client’s hardware resources. In the geographic context, examples for such a server-side analyis and rendering approach are HORA 3D for interactive flooding simulations [14] and applications built on top of ParaView [2]. The disadvantage is that the server needs to be powerful enough to serve all requests at the same time. In addition, interactive applications require a lot of bandwidth for smooth interaction, such as flying over the 3D terrain in Data-Driven Compute Overlays for Interactive Geographic Simulation and Visualization.
@@
The most flexible approach is client-side analysis and rendering: the data to be analyzed remains on the client, and the client handles all interactivity, simulation, and rendering itself. The major disadvantage, clearly, is that the client needs to be powerful enough – in terms of CPU, GPU, and memory – to handle the computation of the overlays and all user interactivity. Client-side overlays can be computed pixel-wise for every frame in the fragment shader based on the raster tiles used for rendering the 3D map. Examples are risk maps computed from local terrain properties [7] or difference maps of flood simulation scenarios [9]. There are some major shortcomings to this approach: First, overlays need to be recomputed whenever user navigation requires a new map tile to be loaded, even if the underlying data does not change. Second, the resolution of the overlay is bound to the resolution of the underlying map tile. Third, pixel-wise analysis is not sufficient for more complex scenarios like avalanche simulation (see Section 4.2). It has been shown that some more complex geospatial analyses can be significantly sped up by GPGPU, using, for instance OpenCL [4] or CUDA [22]. Vollmer et al. [20] compute levels-of-detail of pre-computed static overlay textures on the GPU so that the user can interactively change aggregation parameters. In contrast, we propose a workflow and present an implementation to interactively compute overlay textures directly from the raw data.
@@
Up until recently, graphics APIs for the web were limited to vertex and fragment shaders. With the new web-based graphics and compute GPU API WebGPU, it is now possible to employ compute shaders in the browser. WebGPU is increasingly generating interest for interactive web-based visualization of large data. In particular, it has been used for interactive isosurface extraction [18], graph drawing [5], and aggregate 2D GeoVis [11]. WebGPU is generally seen as a performant successor of WebGL in the context of WebGIS applications [19] and is appreciated for its code portability between web, desktop, and mobile map applications [1]. We extend these works by contributing a workflow, a software architecture, and performance benchmarks for data-driven compute overlays for 3D maps. To the best of our knowledge, weBIGeo is thereby the first online map with interactively computed data-driven overlays.
@@ 2 Data-Driven Compute Overlays
weBIGeo exposes a simple and flexible API to construct compute workflows of varying complexity. Compute workflows are specified as a sequence of compute nodes, similarly to dataflow models for visualization provenance [17] or for steering simulation ensembles [21]. Each compute node has at least one input and one output, as illustrated in the node graph in Fig. 1. Raw input data can come in various forms – from single values, such as the current snow line in the region of interest, over spatial point data stored in a spatial data structure, such as the location of weather stations stored in a kd-tree, to raster data stored in a tile format, such as a digital elevation model (DEM) capturing the ground elevation within a given region. The output of a compute step can be resources like textures, buffers, or vectors. The compute workflow can be configured such that the output of one compute step can serve as input to a subsequent one. This allows arbitrarily reusing and composing compute steps across multiple compute workflow definitions. Compute nodes themselves can represent any CPU or GPU computation. If the input data can be decomposed into small, independent units, a node may execute a compute shader to process each unit in parallel.
@@ 4.2 Avalanches
Gravitational mass flow (GMF) models compute the flow of material under gravity in a given terrain and are commonly used for snow avalanche, landslide, or rock fall simulations. Physically-based GMF models are computationally too expensive for larger-scale simulations and are also very sensitive to input parameters. Larger-scale simulations therefore usually rely on empirical models that use fewer input parameters and can approximate GMF to a certain degree. A state-of-the-art model is FlowPy [3], which is implemented in Python and which we will use for comparison.
As shown in Fig. 1, the compute workflow first fetches elevation tiles of the DEM in the region of interest and computes the normals. By default, we operate on the recommended DEM resolution for avalanche simulations of five to ten meters [10]. The subsequent release point node computes, for each texel of the normal texture in parallel, whether the point is a potential release point of an avalanche. Our release area selection is a very simple approximation, evaluating the local steepness of the terrain as only input parameter, thereby functioning very similarly as the previously described snow cover simulation. For this compute step, the user can control in which texture resolution to evaluate potential release points and which steepness range to consider. More sophisticated models take further parameters – like the upslope area to model the water input or geological maps describing sediment availability – into consideration [10]. Alternatively, we can provide known release points encoded in a pre-computed texture, such as for historic avalanche events, as shown in Data-Driven Compute Overlays for Interactive Geographic Simulation and Visualization. The compute node outputs a binary map indicating whether avalanche particles should be triggered from this point or not.
@@
A comparison between the average model runtimes shows significant speedup of weBIGeo over FlowPy (Section 4.2). On both hardware setups, weBIGeo takes milliseconds for parabola and seconds for vals. In contrast, FlowPy performs in the order of seconds for parabola and hours for vals. These timings exclude file read and write times. As weBIGeo allows to visualize the simulation output directly, file export is only performed on-demand.
Table 1: Model runtime comparison of FlowPy [3] vs weBIGeo, including speedup factor of weBIGeo over FlowPy.
{tabu}
rrrrr Hardware Dataset FlowPy weBIGeo Speedup

### v2

@@ 1 Background: (Web-Based) GeoVis
A node graph specifying the avalanche workflow. Nodes have names, inputs and outputs. The outputs of each node are connected to inputs of subsequent nodes.
@@
Common geographic visualization techniques are glyphs placed on top of maps and color overlays [luo2003visualizing]. The focus of our work lies on overlays. Here, we review three principle paradigms to data-driven map overlays particularly for the web:
@@
Map tiles are the foundation for all popular online map applications, where map clients request pre-rasterized map images depending on the currently shown zoom level and visible field of view [quinn2010predictive] from a dedicated server – also known as “slippy maps” [haklay2008openstreetmap]. Data-driven overlays can be similarly pre-rendered into offline tiles and overlaid onto a base map [egenhofer1993exploratory, reddy1999terravision]. A well-known open-source library to build layered maps is OpenLayers [santiago2015book]. Offline tiles are useful when the data to be overlaid itself is big and static. The disadvantage is that the server has to re-compute the tiles whenever the underlying data changes. Users have no or only very limited possibilities to interact with the data.
@@
An alternative approach is to render the final images to be viewed by the client directly on the server. The client only forwards users’ interaction with the map (e.g., panning and zooming) and with the data (e.g., filtering) to the server and finally shows the image transmitted by the server. The application is therefore largely independent from the client’s hardware resources. In the geographic context, examples for such a server-side analyis and rendering approach are HORA 3D for interactive flooding simulations [rauer2024hora] and applications built on top of ParaView [cedilnik2006remote]. The disadvantage is that the server needs to be powerful enough to serve all requests at the same time. In addition, interactive applications require a lot of bandwidth for smooth interaction, such as flying over the 3D terrain in Data-Driven Compute Overlays for Interactive Geographic Simulation and Visualization.
@@
The most flexible approach is client-side analysis and rendering: the data to be analyzed remains on the client, and the client handles all interactivity, simulation, and rendering itself. The major disadvantage, clearly, is that the client needs to be powerful enough – in terms of CPU, GPU, and memory – to handle the computation of the overlays and all user interactivity. Client-side overlays can be computed pixel-wise for every frame in the fragment shader based on the raster tiles used for rendering the 3D map. Examples are risk maps computed from local terrain properties [eschner-2023-evl] or difference maps of flood simulation scenarios [heitzler2017gpu]. There are some major shortcomings to this approach: First, overlays need to be recomputed whenever user navigation requires a new map tile to be loaded, even if the underlying data does not change. Second, the resolution of the overlay is bound to the resolution of the underlying map tile. Third, pixel-wise analysis is not sufficient for more complex scenarios like avalanche simulation (see Section 4.2). It has been shown that some more complex geospatial analyses can be significantly sped up by GPGPU, using, for instance OpenCL [dubel2017visualizing] or CUDA [xia2011accelerating]. Vollmer et al. [vollmer2018hierarchical] compute levels-of-detail of pre-computed static overlay textures on the GPU so that the user can interactively change aggregation parameters. In contrast, we propose a workflow and present an implementation to interactively compute overlay textures directly from the raw data.
@@
Up until recently, graphics APIs for the web were limited to vertex and fragment shaders. With the new web-based graphics and compute GPU API WebGPU, it is now possible to employ compute shaders in the browser. WebGPU is increasingly generating interest for interactive web-based visualization of large data. In particular, it has been used for interactive isosurface extraction [usher2020interactive], graph drawing [dyken2022graphwagu], and aggregate 2D GeoVis [kimmersdorfer2023webgpu]. WebGPU is generally seen as a performant successor of WebGL in the context of WebGIS applications [usta2024webgpu] and is appreciated for its code portability between web, desktop, and mobile map applications [ammann2022maplibre]. We extend these works by contributing a workflow, a software architecture, and performance benchmarks for data-driven compute overlays for 3D maps. To the best of our knowledge, weBIGeo is thereby the first online map with interactively computed data-driven overlays.
@@ 2 Data-Driven Compute Overlays
weBIGeo exposes a simple and flexible API to construct compute workflows of varying complexity. Compute workflows are specified as a sequence of compute nodes, similarly to dataflow models for visualization provenance [silva2007provenance] or for steering simulation ensembles [waser2011nodes]. Each compute node has at least one input and one output, as illustrated in the node graph in Fig. 1. Raw input data can come in various forms – from single values, such as the current snow line in the region of interest, over spatial point data stored in a spatial data structure, such as the location of weather stations stored in a kd-tree, to raster data stored in a tile format, such as a digital elevation model (DEM) capturing the ground elevation within a given region. The output of a compute step can be resources like textures, buffers, or vectors. The compute workflow can be configured such that the output of one compute step can serve as input to a subsequent one. This allows arbitrarily reusing and composing compute steps across multiple compute workflow definitions. Compute nodes themselves can represent any CPU or GPU computation. If the input data can be decomposed into small, independent units, a node may execute a compute shader to process each unit in parallel.
@@ 4.1 Snow Cover
3D Terrain from above. A rectangular region is covered in snow. Top left: Detailed, no snow on fine terrain features such as trees. Top right: zoomed-out version of left, lower displayed map resolution because the terrain is farther away, snow covers entire area. Bottom left: Detailed. Bottom right: Snow overlay still detailed, despite lower displayed map resolution.
@@ 4.2 Avalanches
Gravitational mass flow (GMF) models compute the flow of material under gravity in a given terrain and are commonly used for snow avalanche, landslide, or rock fall simulations. Physically-based GMF models are computationally too expensive for larger-scale simulations and are also very sensitive to input parameters. Larger-scale simulations therefore usually rely on empirical models that use fewer input parameters and can approximate GMF to a certain degree. A state-of-the-art model is FlowPy [d2022flow], which is implemented in Python and which we will use for comparison.
As shown in Fig. 1, the compute workflow first fetches elevation tiles of the DEM in the region of interest and computes the normals. By default, we operate on the recommended DEM resolution for avalanche simulations of five to ten meters [horton2013flow]. The subsequent release point node computes, for each texel of the normal texture in parallel, whether the point is a potential release point of an avalanche. Our release area selection is a very simple approximation, evaluating the local steepness of the terrain as only input parameter, thereby functioning very similarly as the previously described snow cover simulation. For this compute step, the user can control in which texture resolution to evaluate potential release points and which steepness range to consider. More sophisticated models take further parameters – like the upslope area to model the water input or geological maps describing sediment availability – into consideration [horton2013flow]. Alternatively, we can provide known release points encoded in a pre-computed texture, such as for historic avalanche events, as shown in Data-Driven Compute Overlays for Interactive Geographic Simulation and Visualization. The compute node outputs a binary map indicating whether avalanche particles should be triggered from this point or not.
@@
2D top-down view of our datasets, used for evaluating our avalanche simulation. The release points and computed avalanche path are overlaid.
@@
A comparison between the average model runtimes shows significant speedup of weBIGeo over FlowPy (Table 1). On both hardware setups, weBIGeo takes milliseconds for parabola and seconds for vals. In contrast, FlowPy performs in the order of seconds for parabola and hours for vals. These timings exclude file read and write times. As weBIGeo allows to visualize the simulation output directly, file export is only performed on-demand.
Table 1: Model runtime comparison of FlowPy [d2022flow] vs weBIGeo, including speedup factor of weBIGeo over FlowPy.
Hardware Dataset FlowPy weBIGeo Speedup